OpenStudy (anonymous):

Solve the system of equations by graphing. x+y=11, x=y-7

OpenStudy (anonymous):

Well here's what x+y=11 looks like as a graph http://www.wolframalpha.com/input/?i=+x%2By%3D11 And here's what x=y-7 looks like http://www.wolframalpha.com/input/?i=x%3Dy-7

OpenStudy (anonymous):

Algebra: Linear Equations, Graphs, Slope Solvers Lessons Answers archive Quiz In Depth Click here to see ALL problems on Linear-equations Question 103396: solve the system of equation by graphing x+y=11 x-y=7 (Scroll Down for Answer!) Did you know that Algebra.Com has hundreds of free volunteer tutors who help people with math homework? Anyone can ask a math question, and most questions get answers! OR get immediate PAID help on: Go!!! Answer by Fombitz(13828) (Show Source): You can put this solution on YOUR website! Looks like the lines intersect at x=9. When x = 9, then using the first equation: When x = 9, then using the second equation: So the intersection point (solution) is (9,2).